OpenStudy (anonymous):

what is the simplified form of (15xy^2/x^2-5x+6) divided by (5x^2y/2x^2-7x+3)

OpenStudy (anonymous):

\[ \frac {\frac{15 x y^2}{x^2-5 x+6} } { \frac{5 x^2 y}{2 x^2-7 x+3}}=\frac{15 x y^2}{x^2-5 x+6} \frac{2 x^2-7 x+3}{5 x^2 y} \]

OpenStudy (anonymous):

\[ \frac {\frac{15 x y^2}{x^2-5 x+6} } { \frac{5 x^2 y}{2 x^2-7 x+3}}=\frac{15 x y^2}{x^2-5 x+6} \frac{2 x^2-7 x+3}{5 x^2 y}=\\ \frac{15 x y^2}{(x-3)(x-2)} \frac{(x-3) (2 x-1)}{5 x^2 y} \]

OpenStudy (anonymous):

\[ \frac{15 x y^2}{(x-3)(x-2)} \frac{(x-3) (2 x-1)}{5 x^2 y}= \frac{3 y}{(x-2)} \frac{ (2 x-1)}{ x }=\frac{3y(2x−1)}{x(x-2)} \]

OpenStudy (anonymous):

Please let me know if you need any help with any step.